If the length of the hour and the minute hands of a clock are 4 cm and 6 cm,
respectively, what is the distance in cm between the tips of the hands at two
o'clock?

Answer by lwsshak3(11628)

If the length of the hour and the minute hands of a clock are 4 cm and 6 cm,
respectively, what is the distance in cm between the tips of the hands at two
o'clock?
**
At 2 o'clock, the included angle between the hour and minute hands=60º
let x=distance between the tips of the hands at 2 o'clock
using law of cosines: c^2=a^2+b^2-2*a*b*cosC
x^2=4^2+6^2-2*4*6*cos60º
x^2=16+36-48*1/2
x^2=28
x≈5.29
distance between the tips of the hands at 2 o'clock=5.29 cm